OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi) show all algebra 2sin^2 x-sinx= 1

OpenStudy (raden):

2sin^2 x - sinx = 1 2sin^2 x - sinx - 1 = 0 (2sinx+1)(sinx-1) = 0 for the zeroes, take each factor then setting be zero 2sinx + 1 = 0 or sinx-1=0 for the 2sinx+1=0 2sinx = -1 sinx=-1/2 x = 210 , 330 for the sinx-1=0 sinx = 1 x = 90 sso, the solution are 90,210,330